OpenStudy (anonymous):

Please help????? :33 create a linear equation for this table of values: x y 4 -7.75 3 -5.75 2 -3.75 1 -1.75 0 0.25 -1 2.25

OpenStudy (mathstudent55):

Look at (0, 0.25). That means the equation is y = mx + 0.25 since when x = 0, you get y = m(0) + 0.25 y = 0.25 To get m, just choose two points anf find the slope, m. Then substitute m in y = mx + 0.25 with the value of the slope.